A Complete Bibliography of Publications in Reviews of Modern Physics (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.03

Title word cross-reference

1/f [DH81, Wei88]. ^{+} [Rub85]. ^{−} [Rub85]. ^{0} [Rub85]. ^{3} [CF85, RHD83, SV87, SV88, Vol84]. _{c} [BM88]. _{L} [KLLW81]. a [CP87]. e^{+}e^{−} [Dui82]. Λ [FI86]. N [Bic87, Das87, Yaf82]. pp [BC85, BC85]. R(t) [FI86]. s [MS82]. × [KLLW81].

-Si [CP87].

1987A [Tri88].

61 [JP90].

absorption [LCEK81]. accidents [WAA^{+}85]. accumulation [vdM85]. Acoustics [Cam86, MMO83]. Addendum [ER85a, JP90, KM91]. Adiabatic [ARZ88]. adjustment [CT87]. admissible [DKM81]. admixtures [LKH82b, LKH82a]. adolescence [BR87]. advances [Ric81]. aftermath [Tri83]. alloys [Bic87]. almost [Vol84]. alpha [AAF85, WR80]. alpha-nucleon [AAF85]. American [BPA^{+}87, CGA^{+}81, WAA^{+}85]. analytical [Sch89a]. angular [MS82]. antibunching [Pau82]. antiferromagnetism [Faw88]. antiprotons [vdM85]. anvil [Jay83]. application [CGM85]. Applications [Tam86, Ton87, JR83, JG89]. approach [BM88, Kog83]. Approaches [BSG81]. approximation [CW88]. aspects [Ahl80a, Ahl80b, Das87]. astronomical [Eli82]. Astrophysical [OM88]. astrophysics [Fow84]. asymmetry [Fit81]. Asymptotic [Bur80]. atomic [AK82, Del81, dVDS83]. atoms [BR82, Hoi80, JHY83, Lie81, Lie82, Sie82]. attractors [ER85a, ER85b, Ott81]. automata [Wol83].

barrier [BL80]. beams [SL88, Ste89]. behavior [Gre83]. Bénard [Beh85]. between [CF85, HL84, Pau86]. beyond [Bur80]. birth [BR87]. blastwaves [OM88]. blue [WM89]. body [BRT83, BRT84, Mat86]. bonding [Hof88]. boson [CW88, IT87]. bosons [Rub85]. bound [BYG85]. boundary [SP89]. Breaking [HRW86, Adl82, Adl83]. Breakup [dMK85]. broken [Mic80].

Calculable [BRT83, BRT84]. calculation [Mon87]. capacity [YH86]. Capture [WR80, BHL^{+}82]. Carlo [JR83]. cell [Jay83]. cellular [Wol83]. centrifuge [Whi84a, Whi84b]. chaos [ER85a, ER85b]. chaotic [Ott81]. characteristics [Eli82]. charge [Fit81, Grü88, JR83]. charge-conjugation [Fit81]. charge-density [Grü88]. charged [Mar88]. chemical [Hof88]. chemistry [BD81b, BD81a]. chromium [Faw88]. chromodynamics [Kog83, Mon87]. classical [CTD^{+}80, Ich82, Yaf82]. climate [PO84]. clusters [Sar86]. coal [CGA^{+}81]. coherence [DKTM88]. collision [CZ83, dMK85]. collisions [AK82, Del81, FM84, Lan80a, RKMG85]. Color [Pin86]. Complex [Pin87]. composites [MR81]. compound [HRW86]. compound-nucleus [HRW86]. concepts [BY86]. Conceptual [Wei80]. condensed [PT81, Sie82, Sil80, Wei88]. conducting [HKSS88]. conductors [AS87]. Configurations [Mic80]. Confrontation [CF85]. conjugation [Fit81]. connected [ARZ88]. constant [FI86, JHY83, Wei89]. constants [CT87]. construction [GRS83]. convection [Beh85]. cooling [Ste86, vdM85]. Cosmological [BBF89, FI86, Wei89]. Cosmology [DZ81]. Coulomb [Min87]. coupled [CTD^{+}80, Ich82]. covalent [JR83]. Coverage [Pin87]. CP [Cro81]. Critical [Ahl80c, MR81, FI86, HS89, Wil83]. cross [BC85, RKMG85]. crystal [Hoi80, Lan80b]. Crystalline [WM89]. Crystals [Bri85]. Current [Tho80a, Fis87, Gal83, KLLW81].

Damping [BBB83]. decays [O’D81]. deep [Bur80]. defects [FGP89, Mic80]. degenerate [Ich82]. degrees [MW88]. density [BD81a, BD81b, Faw88, Grü88, Ich82, JG89]. dependence [MS82]. detection [TF85]. detectors [Bic88]. determination [KLLW81]. determined [Hoi80]. deuteron [BF80]. development [BSG81, HBE87, Rus87]. deviations [KLLW81]. device [Gre83]. devices [Key89]. Diamond [Jay83]. dielectric [DKM81]. diffraction [Hoi80, ZGS^{+}88]. diffusion [FGP89]. digital [Key89]. dilute [Bic87]. dimensional [ABSO81, AFS82, Min87, Str88, Viv84, Str89]. dimensions [BKL^{+}86]. Direct [FM84, Owe87]. Direct-photon [FM84]. directed [BPA^{+}87]. discoveries [Rig86]. discovery [Fit81]. Disordered [LR85, AS87]. dissipative [Eck81, LCD^{+}87, LCD^{+}95]. distributions [CZ83, MS82]. dopant [FGP89]. double [BL80, ZGS^{+}88]. double-humped [BL80]. double-slit [ZGS^{+}88]. drive [Fis87]. Duality [Sav80]. ducts [Cam86]. during [Atk85]. dynamic [HL84, Mac89]. dynamical [Eck81, Ott81]. Dynamics [KM89, KM91, LCD^{+}87, ABSO81, Grü88, Neg82, Spo80, dG85, LCD^{+}95]. Dynamo [Ing81].

E0 [LKH82b, LKH82a]. E2 [LKH82b, LKH82a]. Earth [Ing81]. editor [Ynd81]. Editorial [Laz80, Pin86, Pin87]. effect [Adl82, Adl83, AK82, Yen87, vK86]. effects [AS87, BYG85, Hal86, OP85]. Einstein [Adl82, Adl83]. Elastic [DS84, BC85]. electric [JHY83]. electromagnetic [OP85, Wei80]. Electron [CQ81, RKMG85, Sie82, Bah87, BG86, DS84, HG89, HBE87, Ich82, Lan80a, Rus87, Ton87]. electron-molecule [Lan80a]. Electronic [AFS82, Pic89a, Del81, LR85, Sch89a, Pic89b]. electrons [FGK^{+}88]. electroweak [KLLW81]. elementary [DZ81]. elements [Fow84]. elevated [Atk85]. emission [FGK^{+}88, Sar86]. ended [Gal83]. energies [Jon80]. energy [Ahl80a, Ahl80b, BF80, BC85, BPA^{+}87, BR82, DKTM88, FM84, MS82, Sch89a, Sch89b, Ste89]. enhanced [Mos85]. entropy [Mac89]. equations [AP84, Dei84, Spo80]. equilibrium [AP84]. Ergodic [ER85a, ER85b]. Erratum [Adl83, Ahl80a, BD81a, BRT84, BP88, BP89, EHLQ86, LKH82b, LCD^{+}95, Lie82, Pic89b, SV88, Str89, Wu83]. even [LKH82b, LKH82a]. even-even [LKH82b, LKH82a]. events [Tri82]. evolution [Cha84]. Exact [Tha81]. exchange [Jon80]. Excitation [ABSO81, HG89]. excitations [BBB83]. expansion [Bic87]. expansions [Tho80b]. experiment [CF85, Sch89b]. Experimental [BY86, Fow84, Rub85, Ahl80a, Ahl80b, KLLW81, Ric81]. Experiments [Ste89, Bah87, BU88, Gal83, Hoi80]. Extended [LCEK81]. extragalactic [BBR84].

factors [FI86]. facts [BY86]. Fermi [Lie82, Lie81, Vol84]. fermion [Ste84]. few [Mat86]. few-body [Mat86]. Feynman [Sch86]. field [Adl82, Adl83, AP84, Bra85, Ing81, Neg82, RS86, Sav80, Tha81]. field-theoretical [RS86]. fields [Gal85, JHY83]. films [Atk85, Min87, SL88]. fine [LCEK81]. finite [GPY81]. first [Gal83, Sch89b]. fission [BL80]. flow [MP89, MMO83]. flows [BKL^{+}86]. fluctuations [BFF^{+}81, DH81, SGB^{+}89]. fluid [Dei84, MP89]. fluids [Mar88]. flux [AS87]. fluxes [OP85]. force [CTD^{+}80, MW88]. forces [Sal80]. formalism [JG89]. formation [Lan80b]. formulation [GRS83]. forward [BC85]. foundations [IT87, Wei80]. freedom [Bur80, MW88]. frequency [DH81]. Friedmann [FI86]. fuel [CGA^{+}81]. function [DKM81]. functional [JG89]. functions [HG89]. fundamental [CT87, Sal80]. Fundamentals [Sil80]. fusion [Fre82]. future [Tho80a].

galaxies [Sar86]. gamma [BSG81]. gamma-ray [BSG81]. gas [Hoi80, Min87, Whi84a, Whi84b]. gas-surface [Hoi80]. gases [Cam86, Cam87]. Gauge [Sal80, AP84, DV80, Kog83]. gauges [Lei87]. geometrical [DV80]. geometry [DP88]. Geonium [BG86]. Geophysics [STM^{+}87]. glasses [BY86]. gluon [McL86]. graphite [CFG81]. gravitating [SZ89]. Gravitational [Tho80a, Tho80b]. Gravitational-wave [Tho80a]. gravity [Adl82, Adl83, Alv89, STM^{+}87]. greatest [Tri88]. group [CGM85, Gal85, Wil83]. groups [CGM85]. growth [Atk85, Lan80b]. Guest [Laz80]. gyro [CGBP^{+}85].

H [CP87]. hadron [Mon87]. Hall [Yen87, vK86]. Hamiltonian [Spo80, Viv84]. Heat [JP89, JP90]. heating [GM87]. heavily [Ahl80a, Ahl80b]. Heavy [Ste84]. Heavy-fermion [Ste84]. helioseismology [BU88]. helium [Beh85, BBF89, CFG81]. helium-graphite [CFG81]. Hidden [Mic80]. High [BC85, dVDS83, BM88, BF80, DKTM88, FM84, Ich82, Jay83, Jon80, Pic89b, Pic89a, Sch89b, Ste89, Whi84b]. high-density [Ich82]. High-energy [BC85, BF80, DKTM88, FM84, Sch89b, Ste89]. high-pressure [Jay83]. high-speed [Whi84b]. High-spin [dVDS83]. high-T [BM88]. high-temperature [Pic89b, Pic89a]. highly [KKR82]. holography [Ton87]. humped [BL80]. hydrodynamics [Son87]. hydrogen [CP87]. hydrogens [Sil80]. hyperfine [BYG85].

Ideal [Fre82]. II [Cam87, MR81, Tri83, Whi84b]. illustrating [Gre83]. Illustrations [Pin86]. impact [HG89]. imprinting [FS83]. impurities [CQ81]. increasing [Mac89]. independent [Pau86]. inelastic [Bur80]. inflationary [Bra85]. information [YH86]. Instabilities [Lan80b, Viv84]. instantons [GPY81]. integrability [Tha81]. integrable [KM89, KM91]. Integral [Yen87]. interacting [CW88, IT87]. Interaction [Cam87, JHY83, SL88, AAF85, CFG81, Hoi80]. interactions [HL84, Wei80]. interfaces [SL88]. Interference [Pau86]. interferometer [Gre83]. intermediate [Rub85]. intermittency [SZ89]. internal [Cam87]. interpretation [Cra86]. Introduction [Jac80, Lei87, Alv89]. investigations [Jay83]. ion [BG86]. ionized [CQ81]. ionizing [Ahl80a, Ahl80b]. ions [JHY83]. isospin [HRW86]. Issues [CTD^{+}80].

jets [Cam86, Owe87]. Josephson [HL84]. junctions [HL84].

Kepler [Tri88]. Kinetic [Spo80]. kinetics [Wei88]. Kondo [AFL83].

Large [Owe87, Yaf82, Bic87, Das87, SZ89]. large- [Bic87, Das87]. Large-momentum-transfer [Owe87]. large-scale [SZ89]. laser [CGBP^{+}85, Ste86]. lasers [BSG81]. lattice [Kog83, Mat86]. law [STM^{+}87]. leading [Bur80]. Letter [Ynd81]. level [Fan83]. levels [BR82]. light [Sch82]. limitations [LCEK81]. limits [KLLW81, Spo80, Yaf82]. lines [AK82]. liquid [Beh85, Vol84]. liquids [Ich82, WM89]. localized [Vol84]. look [Hof88]. Lorentz [vB82]. loss [Ahl80a, Ahl80b]. Low [DH81, Ahl80c]. Low-frequency [DH81].

M1 [LKH82b, LKH82a]. Magnetic [AS87, Bic87, Cam87, DS84, Fre82, Ing81, JHY83, Tay86]. Magnetism [CF85, RHD83]. magnetohydrodynamic [Fre82]. magnetospheres [Mic82]. many [BRT83, BRT84, Viv84]. many-body [BRT83, BRT84]. many-dimensional [Viv84]. Markovian [Spo80]. masers [Eli82]. masses [Mon87]. Massive [BP87, BP88, BP89]. materials [JR83, MP89]. matrix [BFF^{+}81]. matter [DKM81, KP89, PT81, Sie82, Wei88]. mean [Neg82]. mean-field [Neg82]. measurement [CTD^{+}80, YH86]. Measurements [HG89]. mechanical [CTD^{+}80, HBE87]. mechanics [Cra86, Wol83, Yaf82]. medium [SZ89]. melting [Str88, Str89]. mesons [O’D81]. metal [Hal86, MR81]. metals [GM87, HBE87, RS86]. method [JR83]. methods [AP84, BRT83, BRT84, Bra85, Gal85, RS86]. microscope [Rus87]. microscopy [BR87, Rus87]. microscopy-from [BR87]. millimeter [TF85]. Mills [DV80, Jac80]. minimal [KLLW81]. Mod [JP90]. model [IT87, Wu82, Led89, Wu83]. models [BHL^{+}82, BU88, Bra85, vB82]. Modern [Pin86]. modes [Cam87]. Molecular [CP87, Sil80]. molecule [Lan80a]. molecules [JHY83, Lie81, Sie82, Lie82]. momentum [Owe87]. Monte [JR83]. motion [Dei84]. motions [Ott81]. multiply [ARZ88]. Multipole [Tho80b, LKH82b, LKH82a]. muonic [BR82]. Muonium [Pat88].

nearly [KM89, KM91]. neutral [AK82, Gal83, KLLW81]. neutral-current [Gal83]. Neutrino [Bah87, KP89, BU88, BP87, BP88, BP89, Sch89b, Ste89]. Neutrino-electron [Bah87]. neutrinos [BHL^{+}82, BP87, BP88, BP89, Led89]. neutron [Gre83]. neutrons [WR80, ZGS^{+}88]. NMR [Rig86]. noise [DH81, Wei88]. noncovariant [Lei87]. nonexponential [Wei88]. Nonlinear [Blo82]. nonresonant [AK82]. nonspecialists [Yen87]. Normal [Vol84, GM87, MR81]. nuclear [BBB83, FS83, Fow84, Neg82, WAA^{+}85, dMK85]. nuclei [DS84, LKH82a, SGB^{+}89, dVDS83, LKH82b]. nucleon [AAF85, MW88]. nucleon-nucleon [MW88]. nucleus [HRW86]. Numerical [Mon87].

observation [Rub85]. Observations [Led89]. one [ABSO81]. one-dimensional [ABSO81]. open [BY86]. optical [HG89, YH86]. optics [Blo82]. optoacoustic [PT81]. order [Bur80]. origin [Cro81, Fow84, Mac89]. oscillations [BP87, KP89, Son87, BP88, BP89]. oscillator [CTD^{+}80]. other [Wei88]. oxide [Atk85, Pic89b, Pic89a]. oxides [BM88].

Pairing [SGB^{+}89]. Pairs [Fan83]. paper [JP90]. parity [Fit81]. Part [Cam86, Cam87, Tri82, Tri83, Whi84a, Whi84b]. Particle [Daw83, BD81a, BD81b, Led89]. particles [Ahl80b, DZ81, Hal86, Owe87, WR80, Ahl80a]. pattern [Lan80b]. Penning [BG86]. Perovskite [BM88]. Perovskite-type [BM88]. perturbation [DP88]. PETRA [Dui82]. phase [CZ83, Sil80]. phase-space [CZ83]. phases [WM89]. phenomena [Ahl80c, Wil83, dMK85, dVDS83]. photoacoustic [Tam86]. Photoelectron [MS82]. Photon [Pau82, FM84]. photons [Owe87, Pau86]. Phys [JP90]. Physical [BPA^{+}87, CGA^{+}81, Eli82, WAA^{+}85, CT87, Hoi80, Jay83]. physicists [RTV86]. Physics [BG86, Key89, PO84, BF80, BFF^{+}81, Dui82, EHLQ84, Led89, McL86, Whi84a, Pin86, EHLQ86]. Pion [Jon80]. planning [CGA^{+}81]. plants [WAA^{+}85]. plasma [Gol84, McL86]. plasmas [Daw83, Fis87, Ich82, Tay86]. Point [FGP89]. polymers [HKSS88]. positron [SL88]. positronium [Ric81]. potential [Hoi80, Sch89a]. potential-energy [Sch89a]. Potts [Wu83, Wu82]. power [WAA^{+}85]. precession [Rig86]. predicted [BHL^{+}82]. Preparation [YH86]. pressure [Jay83]. principle [CTD^{+}80]. Principles [Whi84a, Whi84b, GRS83]. Probing [CFG81]. problem [AFL83, Mat86, MP89, Wei89]. Problems [MMO83]. processes [Atk85, Bur80, Sch86, dMK85]. Production [FS83, BBF89, CGA^{+}81, FM84, Owe87, RKMG85]. projectiles [dMK85]. properties [AFS82, KHL^{+}80, Sil80, WCR^{+}84, vB82]. prospects [JG89, Tho80a]. proton [RKMG85]. protons [WR80]. pulsar [Mic82]. Pulsed [PT81].

QCD [DKTM88, GPY81]. QED [BYG85]. Quantized [SV87, vK86, SV88]. Quantum [Alv89, Bra85, CZ83, Hal86, RS86, Rig86, TF85, Adl82, Adl83, ARZ88, CTD^{+}80, Cra86, Gre83, HBE87, Jac80, Kog83, Mon87, OP85, Tha81, YH86, Yen87]. quantum-mechanical [CTD^{+}80, HBE87]. quark [McL86, MW88]. quark-gluon [McL86]. quartz [Bri85]. quasifree [FGK^{+}88]. quasimolecules [Anh85]. quest [Fow84]. questions [BY86].

radiation [Tho80b]. Radiative [O’D81]. radio [BBR84]. radioactive [KKR82]. radionuclide [WAA^{+}85]. Random [BFF^{+}81, ABSO81]. Random-matrix [BFF^{+}81]. rapidly [SGB^{+}89]. rates [BHL^{+}82]. ray [BSG81, LCEK81, Sar86]. Rayleigh [Beh85]. rays [Anh85]. reactions [DKTM88, HRW86, WR80]. real [MP89]. Recoil [BYG85]. reconnection [Tay86, Whi86]. related [Lie81, Lie82]. relativistic [Ahl80a, Ahl80b]. Relaxation [AP84, Tay86]. release [WAA^{+}85]. Renormalization [Gal85, Wil83]. Report [BPA^{+}87, CGA^{+}81, WAA^{+}85]. repositories [KKR82]. representation [CGM85, Sch89a]. research [CGA^{+}81, Ric81, Tho80a]. resistance [SP89]. Resistive [Whi86]. resonators [Bri85]. results [Alv89]. Rev [JP90]. Review [Bic87, Dui82, KHL^{+}80, Whi84a, Whi84b, WCR^{+}84, HS89, KLLW81]. Reviews [Pin86]. Riemann [MP89]. ring [CGBP^{+}85]. Roads [Eck81]. role [BD81a, BD81b]. rotating [SGB^{+}89, Son87]. rotation [Whi84b]. rules [Mar88].

safety [KKR82]. scalar [Gal85]. Scale [FI86, SZ89]. Scanning [BR87]. scattering [Bah87, BRT83, BC85, CQ81, DS84, BRT84]. science [BPA^{+}87]. search [Cro81]. sections [BC85, RKMG85]. Self [GM87, SZ89]. self-gravitating [SZ89]. Self-heating [GM87]. semiclassical [Ste86]. semiconductors [CQ81, JR83, Pat88]. sensing [Tam86]. separation [Whi84a]. setting [DV80]. severe [WAA^{+}85]. Shell [IT87]. Shell-model [IT87]. Si [CP87]. sign [DKM81]. silicon [Bic88, FGP89]. simplified [BBF89]. simulation [Daw83]. since [Tri88]. Single [ZGS^{+}88, BD81a, BD81b, BG86, Hoi80]. Single- [ZGS^{+}88]. single-crystal [Hoi80]. single-particle [BD81a, BD81b]. size [Hal86]. slit [ZGS^{+}88]. slow [Del81, Wei88]. Society [BPA^{+}87, CGA^{+}81, WAA^{+}85]. Solar [BU88, BHL^{+}82, Bah87]. solid [CF85, RHD83, Sil80]. solids [DH81, FS83]. Solitons [HKSS88, KM89, KM91]. Solution [AFL83, JR83]. solutions [Dei84]. Some [Das87, Alv89]. sound [Cam87]. sources [BBR84]. space [CZ83, CGM85, Sch86]. space-time [Sch86]. spectral [AK82]. Spectroscopy [Sch82, Blo82, Mos85, PT81, Sie82]. spectrum [BFF^{+}81]. speed [Whi84b]. Spin [BY86, Faw88, dVDS83]. Spin-density-wave [Faw88]. Spontaneous [FGK^{+}88]. stability [Cha84, Gal85, MR81]. Standard [BHL^{+}82, Led89]. stars [Cha84]. state [LCD^{+}87, LCD^{+}95, MR81]. states [BYG85, Pat88, Rig86, YH86]. Static [HL84, DKM81, Sil80]. statics [dG85]. Statistical [Wol83, Sav80, Tha81]. status [Tho80a]. stimulated [FGK^{+}88]. Stochastic [vdM85, vB82]. Straggling [Bic88]. Strange [Ott81, ER85a, ER85b, Gre83]. strength [BFF^{+}81]. strengths [LCEK81]. string [DP88]. Strong [Gol84]. Strongly [Ich82]. structural [LCEK81]. structure [BYG85, FS83, KLLW81, LCEK81, Neg82, Pic89b, Pic89a, SZ89]. structure-its [LCEK81]. structures [SZ89]. subshells [MS82]. Sum [Mar88]. Supercollider [EHLQ84, EHLQ86]. superconducting [Min87, MR81]. superconducting-normal-metal [MR81]. superconductivity [BM88]. superconductors [GM87, MR81, Pic89a, Pic89b]. superfluid [Min87, SV87, SV88]. superfluid-superconducting [Min87]. superfluids [Son87]. supernova [Tri88]. Supernovae [Tri82, Tri83]. Surface [Mos85, Hoi80]. Surface-enhanced [Mos85]. surfaces [Hof88, Hoi80, Sch89a, SL88]. survey [AAF85]. Symmetry [Mic80, Adl82, Adl83, Cro81, HRW86]. symmetry-breaking [Adl82, Adl83]. synthetic [CGA^{+}81]. system [LCD^{+}87, LCD^{+}95]. Systems [Pin87, ABSO81, AFS82, ARZ88, Eck81, Fan83, Fre82, Gre83, KM89, KM91, LR85, Ott81, Sav80, Ste84, Tha81, Viv84].

T [BM88]. tapestry [Gla80]. Technicolor [Kau83]. techniques [Bic87, Tam86]. technology [BPA^{+}87]. temperature [Ahl80c, Atk85, GPY81, Pic89b, Pic89a]. their [Cha84]. Theoretical [Ahl80b, BY86, Fow84, Hof88, KLLW81, RS86, Ahl80a]. theories [DV80, Das87, Lie81, Lie82]. Theory [BBR84, Del81, Fis87, Mic82, Adl82, Bra85, BG86, CZ83, CF85, DP88, ER85a, ER85b, Fre82, Gal85, Gla80, HBE87, Ing81, Jac80, KLLW81, Kog83, Lan80a, Neg82, RS86, Sav80, Ste86, Tha81, Wei80, Adl83]. Thermal [SP89]. thin [Bic88, SL88]. Thomas [Lie81, Lie82]. Threads [Gla80]. time [Sch86]. times [HS89]. tool [LCEK81]. total [BC85, RKMG85]. tracks [FS83]. transactional [Cra86]. transfer [Owe87]. transitions [Del81, LKH82b, LKH82a]. Transport [Atk85, vB82, ARZ88, JR83, RS86]. trap [BG86]. Tunneling [HS89, BR87]. Turbulence [SZ89, Beh85, Cam86, Eck81, Gol84]. Turbulent [Dei84]. Two [Str88, AFS82, BKL^{+}86, Fan83, Led89, LCD^{+}87, LCD^{+}95, Min87, Rig86, Str89]. Two-dimensional [Str88, AFS82, Min87, Str89]. two-level [Fan83]. two-state [LCD^{+}87, LCD^{+}95]. type [BM88, DV80, MR81]. type-II [MR81].

U [KLLW81]. Ultrametricity [RTV86]. ultraviolet [Gal85]. unbinding [Min87]. uncertainties [BHL^{+}82]. unification [Sal80]. unified [GRS83, Gla80, Wei80]. universe [Bra85, SZ89]. universes [FI86]. until [Whi84a, Whi84b]. use [FS83]. utilization [CGA^{+}81].

values [FI86]. variational [GRS83]. varying [Ing81]. vector [Rub85]. via [Gal85]. violation [Cro81]. violation-the [Cro81]. Viscous [BKL^{+}86]. visualization [Sch86]. Vortex [Son87, Min87]. vortices [SV87, SV88].

W [Rub85]. wastes [KKR82]. wave [Faw88, Tho80a]. wavelengths [TF85]. waves [Cam86, Cam87, Gol84, Grü88, JP89, JP90]. way [Hof88]. Weak [Viv84, CTD^{+}80, KLLW81, Wei80]. weapons [BPA^{+}87]. Wetting [dG85].

X [Anh85, LCEK81, Sar86]. X-ray [LCEK81, Sar86]. X-rays [Anh85].

Yang [DV80, Jac80].

Z [Rub85].

References

Ali:1985:SAN

[AAF85] S. Ali, A. A. Z. Ahmad, and N. Ferdous. A survey of the alpha-nucleon interaction. Reviews of Modern Physics, 57(4): 923–964, October 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.57.923;   http://
rmp.aps.org/abstract/RMP/v57/i4/p923_1.

Alexander:1981:EDR

[ABSO81] S. Alexander, J. Bernasconi, W. R. Schneider, and R. Orbach. Excitation dynamics in random one-dimensional systems. Reviews of Modern Physics, 53(2):175–198, April 1981. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.53.175;
 http://rmp.aps.org/abstract/RMP/v53/i2/p175_1.

Adler:1982:EGS

[Adl82] Stephen L. Adler. Einstein gravity as a symmetry-breaking effect in quantum field theory. Reviews of Modern Physics, 54(3):729–766, July 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.729;   http://rmp.aps.org/abstract/

RMP/v54/i3/p729_1. See erratum [Adl83].

Adler:1983:EEG

[Adl83] Stephen L. Adler. Erratum: Einstein gravity as a symmetrybreaking effect in quantum field theory. Reviews of Modern Physics, 55(3):837, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.837;   http://
rmp.aps.org/abstract/RMP/v55/i3/p837_1. See [Adl82].

Andrei:1983:SKP

[AFL83] N. Andrei, K. Furuya, and J. H. Lowenstein. Solution of the Kondo problem. Reviews of Modern Physics, 55(2):331–402, April 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.331;   http://rmp.aps.org/abstract/
RMP/v55/i2/p331_1.

Ando:1982:EPT

[AFS82] Tsuneya Ando, Alan B. Fowler, and Frank Stern. Electronic properties of two-dimensional systems. Reviews of Modern Physics, 54(2):437–672, April 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.54.437;   http://
rmp.aps.org/abstract/RMP/v54/i2/p437_1.

Ahlen:1980:ETE

[Ahl80a] Steven P. Ahlen. Erratum: Theoretical and experimental aspects of the energy loss of relativistic heavily ionizing particles. Reviews of Modern Physics, 52(3):653, July 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.653;   http://
rmp.aps.org/abstract/RMP/v52/i3/p653_1. See [Ahl80b].

Ahlen:1980:TEA

[Ahl80b] Steven P. Ahlen. Theoretical and experimental aspects of the energy loss of relativistic heavily ionizing particles. Reviews of Modern Physics, 52(1):121–173, January 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.121;
 http://rmp.aps.org/abstract/RMP/v52/i1/p121_1. See erra-

tum [Ahl80a].

Ahlers:1980:CPL

[Ahl80c] Guenter Ahlers. Critical phenomena at low temperature. Reviews of Modern Physics, 52(2):489–503, April 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.489;
 http://rmp.aps.org/abstract/RMP/v52/i2/p489_1.

Allard:1982:ENN

[AK82] Nicole Allard and John Kielkopf. The effect of neutral nonresonant collisions on atomic spectral lines. Reviews of Modern Physics, 54(4):1103–1182, October 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.54.1103;   http:/
/rmp.aps.org/abstract/RMP/v54/i4/p1103_1.

Alvarez:1989:QGI

[Alv89] Enrique Alvarez. Quantum gravity: an introduction to some recent results. Reviews of Modern Physics, 61(3):561–604, July 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.61.561;   http://rmp.aps.org/abstract/
RMP/v61/i3/p561_1.

Anholt:1985:XRQ

[Anh85] R. Anholt. X-rays from quasimolecules. Reviews of Modern Physics, 57(4):995–1053, October 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.995;
 http://rmp.aps.org/abstract/RMP/v57/i4/p995_1.

Adler:1984:RMG

[AP84] Stephen L. Adler and Tsvi Piran. Relaxation methods for gauge field equilibrium equations. Reviews of Modern Physics, 56(1):1–40, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.56.1;   http://rmp.
aps.org/abstract/RMP/v56/i1/p1_1.

Avron:1988:AQT

[ARZ88] J. E. Avron, A. Raveh, and B. Zur. Adiabatic quantum transport in multiply connected systems. Reviews of Modern Physics, 60(4):873–915, October 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.60.873;   http://
rmp.aps.org/abstract/RMP/v60/i4/p873_1.

Aronov:1987:MFE

[AS87] A. G. Aronov and Yu. V. Sharvin. Magnetic flux effects in disordered conductors. Reviews of Modern Physics, 59(3):755–779, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.755;   http://rmp.aps.org/abstract/
RMP/v59/i3/p755_1.

Atkinson:1985:TPD

[Atk85] A. Atkinson. Transport processes during the growth of oxide films at elevated temperature. Reviews of Modern Physics, 57(2):437–470, April 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.57.437;   http://
rmp.aps.org/abstract/RMP/v57/i2/p437_1.

Bahcall:1987:NES

[Bah87] John N. Bahcall. Neutrino-electron scattering and solar neutrino experiments. Reviews of Modern Physics, 59(2):505–521, April 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.505;   http://rmp.aps.org/abstract/
RMP/v59/i2/p505_1.

Bertsch:1983:DNE

[BBB83] G. F. Bertsch, P. F. Bortignon, and R. A. Broglia. Damping of nuclear excitations. Reviews of Modern Physics, 55(1):287–314, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.287;   http://rmp.aps.org/abstract/
RMP/v55/i1/p287_1.

Bernstein:1989:CHP

[BBF89] Jeremy Bernstein, Lowell S. Brown, and Gerald Feinberg. Cosmological helium production simplified. Reviews of Modern Physics, 61(1):25–39, January 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.61.25;   http://rmp.aps.
org/abstract/RMP/v61/i1/p25_1.

Begelman:1984:TER

[BBR84] Mitchell C. Begelman, Roger D. Blandford, and Martin J. Rees. Theory of extragalactic radio sources. Reviews of Modern Physics, 56(2):255–351, April 1984. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.56.255;
 http://rmp.aps.org/abstract/RMP/v56/i2/p255_1.

Block:1985:HEF

[BC85] M. M. Block and R. N. Cahn. High-energy pp̄ and pp forward elastic scattering and total cross sections. Reviews of Modern Physics, 57(2):563–598, April 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.563;
 http://rmp.aps.org/abstract/RMP/v57/i2/p563_1.

Bamzai:1981:ERS

[BD81a] Anjuli S. Bamzai and B. M. Deb. Erratum: The role of single-particle density in chemistry. Reviews of Modern Physics, 53(3):593, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.53.593;   http://
rmp.aps.org/abstract/RMP/v53/i3/p593_1. See [BD81b].

Bamzai:1981:RSP

[BD81b] Anjuli S. Bamzai and B. M. Deb. The role of single-particle density in chemistry. Reviews of Modern Physics, 53(1):95–126, January 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.53.95;   http://rmp.aps.org/
abstract/RMP/v53/i1/p95_1. See erratum [BD81a].

Behringer:1985:RBC

[Beh85] R. P. Behringer. Rayleigh–Bénard convection and turbulence in liquid helium. Reviews of Modern Physics, 57(3):657–687, July 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.657;   http://rmp.aps.org/abstract/
RMP/v57/i3/p657_1.

Bergstrom:1980:DHE

[BF80] Lars Bergström and Sverker Fredriksson. The deuteron in highenergy physics. Reviews of Modern Physics, 52(4):675–697, October 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.675;   http://rmp.aps.org/abstract/
RMP/v52/i4/p675_1.

Brody:1981:RMP

[BFF^{+}81] T. A. Brody, J. Flores, J. B. French, P. A. Mello, A. Pandey, and S. S. M. Wong. Random-matrix physics: spectrum and strength fluctuations. Reviews of Modern Physics, 53(3):385–479, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.385;   http://rmp.aps.org/abstract/
RMP/v53/i3/p385_1.

Brown:1986:GTP

[BG86] Lowell S. Brown and Gerald Gabrielse. Geonium theory: Physics of a single electron or ion in a Penning trap. Reviews of Modern Physics, 58(1):233–311, January 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.233;
 http://rmp.aps.org/abstract/RMP/v58/i1/p233_1.

Bahcall:1982:SSM

[BHL^{+}82] John N. Bahcall, Walter F. Huebner, Stephen H. Lubow, Peter D. Parker, and Roger K. Ulrich. Standard solar models and the uncertainties in predicted capture rates of solar neutrinos. Reviews of Modern Physics, 54(3):767–799, July 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.767;
 http://rmp.aps.org/abstract/RMP/v54/i3/p767_1.

Bickers:1987:RTL

[Bic87] N. E. Bickers. Review of techniques in the large-N expansion for dilute magnetic alloys. Reviews of Modern Physics, 59 (4):845–939, October 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.59.845;   http://
rmp.aps.org/abstract/RMP/v59/i4/p845_1.

Bichsel:1988:STS

[Bic88] Hans Bichsel. Straggling in thin silicon detectors. Reviews of Modern Physics, 60(3):663–699, July 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.663;
 http://rmp.aps.org/abstract/RMP/v60/i3/p663_1.

Bensimon:1986:VFT

[BKL^{+}86] David Bensimon, Leo P. Kadanoff, Shoudan Liang, Boris I. Shraiman, and Chao Tang. Viscous flows in two dimensions. Reviews of Modern Physics, 58(4):977–999, October 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
58.977;   http://rmp.aps.org/abstract/RMP/v58/i4/p977_
1.

Bjornholm:1980:DHF

[BL80] S. Bjørnholm and J. E. Lynn. The double-humped fission barrier. Reviews of Modern Physics, 52(4):725–931, October 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
52.725;   http://rmp.aps.org/abstract/RMP/v52/i4/p725_

Bloembergen:1982:NOS

[Blo82] Nicolaas Bloembergen. Nonlinear optics and spectroscopy. Reviews of Modern Physics, 54(3):685–695, July 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.685;
 http://rmp.aps.org/abstract/RMP/v54/i3/p685_1.

Bednorz:1988:PTO

[BM88] J. Georg Bednorz and K. Alex Müller. Perovskite-type oxides — the new approach to high-T_{c} superconductivity. Reviews of Modern Physics, 60(3):585–600, July 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.585;
 http://rmp.aps.org/abstract/RMP/v60/i3/p585_1.

Bilenky:1987:MNN

[BP87] S. M. Bilenky and S. T. Petcov. Massive neutrinos and neutrino oscillations. Reviews of Modern Physics, 59(3):671–754, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.671;   http://rmp.aps.org/abstract/

RMP/v59/i3/p671_1. See errata [BP88, BP89].

Bilenky:1988:EMN

[BP88] S. M. Bilenky and S. T. Petcov. Erratum: Massive neutrinos and neutrino oscillations. Reviews of Modern Physics, 60(2):575, April 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.575;   http://rmp.aps.org/abstract/
RMP/v60/i2/p575_1. See [BP87, BP89].

Bilenky:1989:EMN

[BP89] S. M. Bilenky and S. T. Petcov. Erratum: Massive neutrinos and neutrino oscillations. Reviews of Modern Physics, 61(1):169, January 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.61.169;   http://rmp.aps.org/abstract/
RMP/v61/i1/p169_1. See [BP87, BP88].

Bloembergen:1987:RAP

[BPA^{+}87] N. Bloembergen, C. K. N. Patel, P. Avizonis, R. G. Clem, A. Hertzberg, T. H. Johnson, T. Marshall, R. B. Miller, W. E.

Morrow, E. E. Salpeter, A. M. Sessler, J. D. Sullivan, J. C. Wyant, A. Yariv, R. N. Zare, A. J. Glass, L. C. Hebel, G. E. Pake, M. M. May, W. K. Panofsky, A. L. Schawlow, C. H. Townes, and H. York (APS Study Group Participants and APS Council Review Committee). Report to The American Physical Society of the study group on science and technology of directed energy weapons. Reviews of Modern Physics, 59(3):S1–S201, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
59.S1;   http://rmp.aps.org/abstract/RMP/v59/i3/pS1_1.

Borie:1982:ELM

[BR82] E. Borie and G. A. Rinker. The energy levels of muonic atoms. Reviews of Modern Physics, 54(1):67–118, January 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
54.67;   http://rmp.aps.org/abstract/RMP/v54/i1/p67_1.

Binnig:1987:STM

[BR87] Gerd Binnig and Heinrich Rohrer. Scanning tunneling microscopyfrom birth to adolescence. Reviews of Modern Physics, 59 (3):615–625, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.59.615;   http://
rmp.aps.org/abstract/RMP/v59/i3/p615_1.

Brandenberger:1985:QFT

[Bra85] Robert H. Brandenberger. Quantum field theory methods and inflationary universe models. Reviews of Modern Physics, 57(1):1–60, January 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.57.1;   http://rmp.aps.org/abstract/
RMP/v57/i1/p1_1.

Brice:1985:CQR

[Bri85] J. C. Brice. Crystals for quartz resonators. Reviews of Modern Physics, 57(1):105–146, January 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.105;
 http://rmp.aps.org/abstract/RMP/v57/i1/p105_1.

Barrett:1983:CMM

[BRT83] R. F. Barrett, B. A. Robson, and W. Tobocman. Calculable methods for many-body scattering. Reviews of Modern Physics, 55(1):155–243, January 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.155;
 http://rmp.aps.org/abstract/RMP/v55/i1/p155_1. See erra-

tum [BRT84].

Barrett:1984:ECM

[BRT84] R. F. Barrett, B. A. Robson, and W. Tobocman. Erratum: Calculable methods for many-body scattering. Reviews of Modern Physics, 56(3):567, July 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.56.567;   http://
rmp.aps.org/abstract/RMP/v56/i3/p567_1. See [BRT83].

Baldwin:1981:ADG

[BSG81] George C. Baldwin, Johndale C. Solem, and Vitalii I. Gol’danskii. Approaches to the development of gamma-ray lasers. Reviews of Modern Physics, 53(4):687–744, October 1981. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.53.687;
 http://rmp.aps.org/abstract/RMP/v53/i4/p687_1.

Bahcall:1988:SMN

[BU88] John N. Bahcall and Roger K. Ulrich. Solar models, neutrino experiments, and helioseismology. Reviews of Modern Physics, 60(2):297–372, April 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.60.297;   http://
rmp.aps.org/abstract/RMP/v60/i2/p297_1.

Buras:1980:AFD

[Bur80] Andrzej J. Buras. Asymptotic freedom in deep inelastic processes in the leading order and beyond. Reviews of Modern Physics, 52(1):199–276, January 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.199;   http://
rmp.aps.org/abstract/RMP/v52/i1/p199_1.

Binder:1986:SGE

[BY86] K. Binder and A. P. Young. Spin glasses: Experimental facts, theoretical concepts, and open questions. Reviews of Modern Physics, 58(4):801–976, October 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.801;
 http://rmp.aps.org/abstract/RMP/v58/i4/p801_1.

Bodwin:1985:REH

[BYG85] G. T. Bodwin, D. R. Yennie, and M. A. Gregorio. Recoil effects in the hyperfine structure of QED bound states. Reviews of Modern Physics, 57(3):723–782, July 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.723;
 http://rmp.aps.org/abstract/RMP/v57/i3/p723_1.

Campos:1986:WGP

[Cam86] L. M. B. C. Campos. On waves in gases. Part I: Acoustics of jets, turbulence, and ducts. Reviews of Modern Physics, 58 (1):117–182, January 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.58.117;   http://
rmp.aps.org/abstract/RMP/v58/i1/p117_1.

Campos:1987:WGP

[Cam87] L. M. B. C. Campos. On waves in gases. Part II: Interaction of sound with magnetic and internal modes. Reviews of Modern Physics, 59(2):363–463, April 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.363;
 http://rmp.aps.org/abstract/RMP/v59/i2/p363_1.

Cross:1985:MSC

[CF85] M. C. Cross and Daniel S. Fisher. Magnetism in solid ^{3}He: Confrontation between theory and experiment. Reviews of Modern Physics, 57(4):881–921, October 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.881;
 http://rmp.aps.org/abstract/RMP/v57/i4/p881_1.

Cole:1981:PHG

[CFG81] Milton W. Cole, D. R. Frankl, and David L. Goodstein. Probing the helium-graphite interaction. Reviews of Modern Physics, 53(2):199–210, April 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.53.199;   http://
rmp.aps.org/abstract/RMP/v53/i2/p199_1.

Cooper:1981:RAP

[CGA^{+}81] Bernard R. Cooper, Wayne R. Gruner, Larry Anderson, Robert H. Davis, Paul Engelking, Jose D. Garcia, Edward Gerjuoy, Robert I. Jaffee, Philip G. Kosky, Leonidas Petrakis, Robert T. Poe, Richard Pollina, C. Dwight Prater, Robert L. Thomas, Sandor Trajmar, G. J. Dienes, Nancy M. O’Fallon, Harvey Brooks, William A. Fowler, Everett Gorin, Norman F. Ramsey, and M. D. Schlesinger (APS Study Group Participants and Ex Officio Panel Members and APS Council Review Committee). Report to the American Physical Society by the study group on research planning for coal utilization and synthetic fuel production. Reviews of Modern Physics, 53(4):S1–S168, October 1981. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
53.S1;   http://rmp.aps.org/abstract/RMP/v53/i4/pS1_1.

Chow:1985:RLG

[CGBP^{+}85] W. W. Chow, J. Gea-Banacloche, L. M. Pedrotti, V. E. Sanders, W. Schleich, and M. O. Scully. The ring laser gyro. Reviews of Modern Physics, 57(1):61–104, January 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
57.61;   http://rmp.aps.org/abstract/RMP/v57/i1/p61_1.

Chen:1985:RGA

[CGM85] Jin-Quan Chen, Mei-Juan Gao, and Guang-Qun Ma. The representation group and its application to space groups. Reviews of Modern Physics, 57(1):211–278, January 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.211;
 http://rmp.aps.org/abstract/RMP/v57/i1/p211_1.

Chandrasekhar:1984:STE

[Cha84] S. Chandrasekhar. On stars, their evolution and their stability. Reviews of Modern Physics, 56(2):137–147, April 1984. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.56.137;
 http://rmp.aps.org/abstract/RMP/v56/i2/p137_1.

Chabal:1987:MHS

[CP87] Y. J. Chabal and C. K. N. Patel. Molecular hydrogen in a-Si:H. Reviews of Modern Physics, 59(4):835–844, October 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
59.835;   http://rmp.aps.org/abstract/RMP/v59/i4/p835_
1.

Chattopadhyay:1981:ESI

[CQ81] D. Chattopadhyay and H. J. Queisser. Electron scattering by ionized impurities in semiconductors. Reviews of Modern Physics, 53(4):745–768, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.53.745;   http://
rmp.aps.org/abstract/RMP/v53/i4/p745_1.

Cramer:1986:TIQ

[Cra86] John G. Cramer. The transactional interpretation of quantum mechanics. Reviews of Modern Physics, 58(3):647–687, July 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.647;   http://rmp.aps.org/abstract/
RMP/v58/i3/p647_1;   https://en.wikipedia.org/wiki/Transactional_
interpretation.

Cronin:1981:CSV

[Cro81] James W. Cronin. CP symmetry violation-the search for its origin. Reviews of Modern Physics, 53(3):373–383, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
53.373;   http://rmp.aps.org/abstract/RMP/v53/i3/p373_
1.

Cohen:1987:AFP

[CT87] E. Richard Cohen and Barry N. Taylor. The 1986 adjustment of the fundamental physical constants. Reviews of Modern Physics, 59(4):1121–1148, October 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.59.1121;   http://
rmp.aps.org/abstract/RMP/v59/i4/p1121_1.

Caves:1980:MWC

[CTD^{+}80] Carlton M. Caves, Kip S. Thorne, Ronald W. P. Drever, Vernon D. Sandberg, and Mark Zimmermann. On the measurement of a weak classical force coupled to a quantum-mechanical oscillator. I. Issues of principle. Reviews of Modern Physics, 52(2):341–392, April 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.341;   http://
rmp.aps.org/abstract/RMP/v52/i2/p341_1.

Casten:1988:IBA

[CW88] R. F. Casten and D. D. Warner. The interacting boson approximation. Reviews of Modern Physics, 60(2):389–469, April 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.389;   http://rmp.aps.org/abstract/
RMP/v60/i2/p389_1.

Carruthers:1983:QCT

[CZ83] P. Carruthers and F. Zachariasen. Quantum collision theory with phase-space distributions. Reviews of Modern Physics, 55 (1):245–285, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.245;   http://
rmp.aps.org/abstract/RMP/v55/i1/p245_1.

Das:1987:SAL

[Das87] Sumit R. Das. Some aspects of large-N theories. Reviews of Modern Physics, 59(1):235–261, January 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.235;
 http://rmp.aps.org/abstract/RMP/v59/i1/p235_1.

Dawson:1983:PSP

[Daw83] John M. Dawson. Particle simulation of plasmas. Reviews of Modern Physics, 55(2):403–447, April 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.403;
 http://rmp.aps.org/abstract/RMP/v55/i2/p403_1.

Deissler:1984:TSE

[Dei84] R. G. Deissler. Turbulent solutions of the equations of fluid motion. Reviews of Modern Physics, 56(2):223–254, April 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
56.223;   http://rmp.aps.org/abstract/RMP/v56/i2/p223_
1.

Delos:1981:TET

[Del81] John B. Delos. Theory of electronic transitions in slow atomic collisions. Reviews of Modern Physics, 53(2):287–357, April 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.287;   http://rmp.aps.org/abstract/
RMP/v53/i2/p287_1.

deGennes:1985:WSD

[dG85] P. G. de Gennes. Wetting: statics and dynamics. Reviews of Modern Physics, 57(3):827–863, July 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.827;
 http://rmp.aps.org/abstract/RMP/v57/i3/p827_1.

Dutta:1981:LFF

[DH81] P. Dutta and P. M. Horn. Low-frequency fluctuations in solids: 1/f noise. Reviews of Modern Physics, 53(3):497–516, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.497;   http://rmp.aps.org/abstract/
RMP/v53/i3/p497_1.

Dolgov:1981:ASS

[DKM81] O. V. Dolgov, D. A. Kirzhnits, and E. G. Maksimov. On an admissible sign of the static dielectric function of matter. Reviews of Modern Physics, 53(1):81–93, January 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
53.81;   http://rmp.aps.org/abstract/RMP/v53/i1/p81_1.

Dokshitzer:1988:QCH

[DKTM88] Yu. L. Dokshitzer, V. A. Khoze, S. I. Troyan, and A. H. Mueller. QCD coherence in high-energy reactions. Reviews of Modern Physics, 60(2):373–388, April 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.373;
 http://rmp.aps.org/abstract/RMP/v60/i2/p373_1.

deMeijer:1985:BPN

[dMK85] R. J. de Meijer and R. Kamermans. Breakup phenomena in nuclear collision processes with He projectiles. Reviews of Modern Physics, 57(1):147–209, January 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.147;
 http://rmp.aps.org/abstract/RMP/v57/i1/p147_1.

DHoker:1988:GSP

[DP88] Eric D’Hoker and D. H. Phong. The geometry of string perturbation theory. Reviews of Modern Physics, 60(4):917–1065, October 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.917;   http://rmp.aps.org/abstract/
RMP/v60/i4/p917_1.

Donnelly:1984:EME

[DS84] T. William Donnelly and Ingo Sick. Elastic magnetic electron scattering from nuclei. Reviews of Modern Physics, 56(3):461–566, July 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.56.461;   http://rmp.aps.org/abstract/
RMP/v56/i3/p461_1.

Duinker:1982:RPP

[Dui82] P. Duinker. Review of e^{+}e^{−} physics at PETRA. Reviews of Modern Physics, 54(2):325–387, April 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.325;
 http://rmp.aps.org/abstract/RMP/v54/i2/p325_1.

Daniel:1980:GSG

[DV80] M. Daniel and C. M. Viallet. The geometrical setting of gauge theories of the Yang–Mills type. Reviews of Modern Physics, 52(1):175–197, January 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.175;   http://
rmp.aps.org/abstract/RMP/v52/i1/p175_1.

deVoigt:1983:HSP

[dVDS83] M. J. A. de Voigt, J. Dudek, and Z. Szymański. High-spin phenomena in atomic nuclei. Reviews of Modern Physics, 55 (4):949–1046, October 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.949;   http://
rmp.aps.org/abstract/RMP/v55/i4/p949_1.

Dolgov:1981:CEP

[DZ81] A. D. Dolgov and Ya. B. Zeldovich. Cosmology and elementary particles. Reviews of Modern Physics, 53(1):1–41, January 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.53.1;   http://rmp.aps.org/abstract/RMP/v53/
i1/p1_1.

Eckmann:1981:RTD

[Eck81] J.-P. Eckmann. Roads to turbulence in dissipative dynamical systems. Reviews of Modern Physics, 53(4):643–654, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.643;   http://rmp.aps.org/abstract/
RMP/v53/i4/p643_1.

Eichten:1984:SP

[EHLQ84] E. Eichten, I. Hinchliffe, K. Lane, and C. Quigg. Supercollider physics. Reviews of Modern Physics, 56(4):579–707, October 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.56.579;   http://rmp.aps.org/abstract/

RMP/v56/i4/p579_1. See erratum [EHLQ86].

Eichten:1986:ESP

[EHLQ86] E. Eichten, I. Hinchliffe, K. Lane, and C. Quigg. Erratum: Supercollider physics. Reviews of Modern Physics, 58(4):1065–1073, October 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.1065;   http://rmp.aps.org/abstract/
RMP/v58/i4/p1065_1. See [EHLQ84].

Elitzur:1982:PCA

[Eli82] Moshe Elitzur. Physical characteristics of astronomical masers. Reviews of Modern Physics, 54(4):1225–1260, October 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.54.1225;   http://rmp.aps.org/abstract/RMP/
v54/i4/p1225_1.

Eckmann:1985:AET

[ER85a] J.-P. Eckmann and D. Ruelle. Addendum: Ergodic theory of chaos and strange attractors. Reviews of Modern Physics, 57(4):1115, October 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.1115;   http://rmp.aps.org/abstract/
RMP/v57/i4/p1115_1. See [ER85b].

Eckmann:1985:ETC

[ER85b] J.-P. Eckmann and D. Ruelle. Ergodic theory of chaos and strange attractors. Reviews of Modern Physics, 57(3):617–656, July 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.617;   http://rmp.aps.org/abstract/

RMP/v57/i3/p617_1. See addendum [ER85a].

Fano:1983:PTL

[Fan83] U. Fano. Pairs of two-level systems. Reviews of Modern Physics, 55(4):855–874, October 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.855;   http://
rmp.aps.org/abstract/RMP/v55/i4/p855_1.

Fawcett:1988:SDW

[Faw88] Eric Fawcett. Spin-density-wave antiferromagnetism in chromium. Reviews of Modern Physics, 60(1):209–283, January 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
60.209;   http://rmp.aps.org/abstract/RMP/v60/i1/p209_
1.

Friedman:1988:SSE

[FGK^{+}88] A. Friedman, A. Gover, G. Kurizki, S. Ruschin, and A. Yariv. Spontaneous and stimulated emission from quasifree electrons. Reviews of Modern Physics, 60(2):471–535, April 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.471;
 http://rmp.aps.org/abstract/RMP/v60/i2/p471_1.

Fahey:1989:PDD

[FGP89] P. M. Fahey, P. B. Griffin, and J. D. Plummer. Point defects and dopant diffusion in silicon. Reviews of Modern Physics, 61(2):289–384, April 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.289;   http://
rmp.aps.org/abstract/RMP/v61/i2/p289_1.

Felten:1986:SFC

[FI86] James E. Felten and Richard Isaacman. Scale factors R(t) and critical values of the cosmological constant Λ in Friedmann universes. Reviews of Modern Physics, 58(3):689–698, July 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
58.689;   http://rmp.aps.org/abstract/RMP/v58/i3/p689_
1.

Fisch:1987:TCD

[Fis87] Nathaniel J. Fisch. Theory of current drive in plasmas. Reviews of Modern Physics, 59(1):175–234, January 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.175;
 http://rmp.aps.org/abstract/RMP/v59/i1/p175_1.

Fitch:1981:DCC

[Fit81] Val L. Fitch. The discovery of charge-conjugation parity asymmetry. Reviews of Modern Physics, 53(3):367–371, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
53.367;   http://rmp.aps.org/abstract/RMP/v53/i3/p367_
1.

Ferbel:1984:DPP

[FM84] T. Ferbel and W. R. Molzon. Direct-photon production in highenergy collisions. Reviews of Modern Physics, 56(2):181–221, April 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.56.181;   http://rmp.aps.org/abstract/
RMP/v56/i2/p181_1.

Fowler:1984:ETN

[Fow84] William A. Fowler. Experimental and theoretical nuclear astrophysics: the quest for the origin of the elements. Reviews of Modern Physics, 56(2):149–179, April 1984. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.56.149;
 http://rmp.aps.org/abstract/RMP/v56/i2/p149_1.

Freidberg:1982:IMT

[Fre82] J. P. Freidberg. Ideal magnetohydrodynamic theory of magnetic fusion systems. Reviews of Modern Physics, 54(3):801–902, July 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.801;   http://rmp.aps.org/abstract/
RMP/v54/i3/p801_1.

Fischer:1983:PUN

[FS83] Bernd Eberhard Fischer and Reimar Spohr. Production and use of nuclear tracks: imprinting structure on solids. Reviews of Modern Physics, 55(4):907–948, October 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.907;
 http://rmp.aps.org/abstract/RMP/v55/i4/p907_1.

Galison:1983:HFN

[Gal83] Peter Galison. How the first neutral-current experiments ended. Reviews of Modern Physics, 55(2):477–509, April 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
55.477;   http://rmp.aps.org/abstract/RMP/v55/i2/p477_
1.

Gallavotti:1985:RTU

[Gal85] Giovanni Gallavotti. Renormalization theory and ultraviolet stability for scalar fields via renormalization group methods. Reviews of Modern Physics, 57(2):471–562, April 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.471;
 http://rmp.aps.org/abstract/RMP/v57/i2/p471_1.

Glashow:1980:TUT

[Gla80] Sheldon Lee Glashow. Towards a unified theory: Threads in a tapestry. Reviews of Modern Physics, 52(3):539–543, July 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.539;   http://rmp.aps.org/abstract/
RMP/v52/i3/p539_1.

Gurevich:1987:SHN

[GM87] A. Vl. Gurevich and R. G. Mints. Self-heating in normal metals and superconductors. Reviews of Modern Physics, 59(4):941–999, October 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.941;   http://rmp.aps.org/abstract/
RMP/v59/i4/p941_1.

Goldman:1984:STP

[Gol84] Martin V. Goldman. Strong turbulence of plasma waves. Reviews of Modern Physics, 56(4):709–735, October 1984. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.56.709;
 http://rmp.aps.org/abstract/RMP/v56/i4/p709_1.

Gross:1981:QIF

[GPY81] David J. Gross, Robert D. Pisarski, and Laurence G. Yaffe. QCD and instantons at finite temperature. Reviews of Modern Physics, 53(1):43–80, January 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.53.43;   http://rmp.aps.
org/abstract/RMP/v53/i1/p43_1.

Greenberger:1983:NID

[Gre83] Daniel M. Greenberger. The neutron interferometer as a device for illustrating the strange behavior of quantum systems. Reviews of Modern Physics, 55(4):875–905, October 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.875;
 http://rmp.aps.org/abstract/RMP/v55/i4/p875_1.

Gerjuoy:1983:UFC

[GRS83] E. Gerjuoy, A. R. P. Rau, and Larry Spruch. A unified formulation of the construction of variational principles. Reviews of Modern Physics, 55(3):725–774, July 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.725;
 http://rmp.aps.org/abstract/RMP/v55/i3/p725_1.

Gruner:1988:DCD

[Grü88] G. Grüner. The dynamics of charge-density waves. Reviews of Modern Physics, 60(4):1129–1181, October 1988. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.60.1129;   http://rmp.aps.org/abstract/RMP/
v60/i4/p1129_1.

Halperin:1986:QSE

[Hal86] W. P. Halperin. Quantum size effects in metal particles. Reviews of Modern Physics, 58(3):533–606, July 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.533;
 http://rmp.aps.org/abstract/RMP/v58/i3/p533_1.

Hoddeson:1987:DQM

[HBE87] Lillian Hoddeson, Gordon Baym, and Michael Eckert. The development of the quantum-mechanical electron theory of metals: 1928–1933. Reviews of Modern Physics, 59(1):287–327, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.287;   http://rmp.aps.org/abstract/
RMP/v59/i1/p287_1.

Heddle:1989:MEI

[HG89] D. W. O. Heddle and Jean W. Gallagher. Measurements of electron impact optical excitation functions. Reviews of Modern Physics, 61(2):221–278, April 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.221;
 http://rmp.aps.org/abstract/RMP/v61/i2/p221_1.

Heeger:1988:SCP

[HKSS88] A. J. Heeger, S. Kivelson, J. R. Schrieffer, and W.-P. Su. Solitons in conducting polymers. Reviews of Modern Physics, 60(3):781–850, July 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.781;   http://rmp.aps.org/abstract/
RMP/v60/i3/p781_1.

Hansen:1984:SDI

[HL84] J. Bindslev Hansen and P. E. Lindelof. Static and dynamic interactions between Josephson junctions. Reviews of Modern Physics, 56(3):431–459, July 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.56.431;   http://
rmp.aps.org/abstract/RMP/v56/i3/p431_1.

Hoffmann:1988:CTW

[Hof88] Roald Hoffmann. A chemical and theoretical way to look at bonding on surfaces. Reviews of Modern Physics, 60(3):601–628, July 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.601;   http://rmp.aps.org/abstract/
RMP/v60/i3/p601_1.

Hoinkes:1980:PIP

[Hoi80] H. Hoinkes. The physical interaction potential of gas atoms with single-crystal surfaces, determined from gas-surface diffraction experiments. Reviews of Modern Physics, 52(4):933–970, October 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.933;   http://rmp.aps.org/abstract/
RMP/v52/i4/p933_1.

Harney:1986:BIS

[HRW86] H. L. Harney, A. Richter, and H. A. Weidenmüller. Breaking of isospin symmetry in compound-nucleus reactions. Reviews of Modern Physics, 58(3):607–645, July 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.607;
 http://rmp.aps.org/abstract/RMP/v58/i3/p607_1.

Hauge:1989:TTC

[HS89] E. H. Hauge and J. A. Støvneng. Tunneling times: a critical review. Reviews of Modern Physics, 61(4):917–936, October 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
61.917;   http://rmp.aps.org/abstract/RMP/v61/i4/p917_
1.

Ichimaru:1982:SCP

[Ich82] Setsuo Ichimaru. Strongly coupled plasmas: high-density classical plasmas and degenerate electron liquids. Reviews of Modern Physics, 54(4):1017–1059, October 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.54.1017;   http:/
/rmp.aps.org/abstract/RMP/v54/i4/p1017_1.

Inglis:1981:DTE

[Ing81] David Rittenhouse Inglis. Dynamo theory of the Earth’s varying magnetic field. Reviews of Modern Physics, 53(3):481–496, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.481;   http://rmp.aps.org/abstract/
RMP/v53/i3/p481_1.

Iachello:1987:SMF

[IT87] F. Iachello and I. Talmi. Shell-model foundations of the interacting boson model. Reviews of Modern Physics, 59(2):339–361, April 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.339;   http://rmp.aps.org/abstract/
RMP/v59/i2/p339_1.

Jackiw:1980:IYM

[Jac80] R. Jackiw. Introduction to the Yang–Mills quantum theory. Reviews of Modern Physics, 52(4):661–673, October 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
52.661;   http://rmp.aps.org/abstract/RMP/v52/i4/p661_
1.

Jayaraman:1983:DAC

[Jay83] A. Jayaraman. Diamond anvil cell and high-pressure physical investigations. Reviews of Modern Physics, 55(1):65–108, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.55.65;   http://rmp.aps.org/
abstract/RMP/v55/i1/p65_1.

Jones:1989:DFF

[JG89] R. O. Jones and O. Gunnarsson. The density functional formalism, its applications and prospects. Reviews of Modern Physics, 61(3):689–746, July 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.689;   http://
rmp.aps.org/abstract/RMP/v61/i3/p689_1.

Johnson:1983:IAM

[JHY83] Bruce R. Johnson, Joseph O. Hirschfelder, and Kuo-Ho Yang. Interaction of atoms, molecules, and ions with constant electric and magnetic fields. Reviews of Modern Physics, 55(1):109–153, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.109;   http://rmp.aps.org/abstract/
RMP/v55/i1/p109_1.

Jones:1980:PEH

[Jon80] Lorella M. Jones. Pion exchange at high energies. Reviews of Modern Physics, 52(3):545–567, July 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.545;
 http://rmp.aps.org/abstract/RMP/v52/i3/p545_1.

Joseph:1989:HW

[JP89] D. D. Joseph and Luigi Preziosi. Heat waves. Reviews of Modern Physics, 61(1):41–73, January 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.41;
 http://rmp.aps.org/abstract/RMP/v61/i1/p41_1. See adden-

dum [JP90].

Joseph:1990:APH

[JP90] D. D. Joseph and Luigi Preziosi. Addendum to the paper “Heat waves” [Rev. Mod. Phys. 61, 41 (1989)]. Reviews of Modern Physics, 62(2):375–391, April 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.62.375;   http://
rmp.aps.org/abstract/RMP/v62/i2/p375_1. See [JP89].

Jacoboni:1983:MCM

[JR83] Carlo Jacoboni and Lino Reggiani. The Monte Carlo method for the solution of charge transport in semiconductors with applications to covalent materials. Reviews of Modern Physics, 55(3):645–705, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.645;   http://
rmp.aps.org/abstract/RMP/v55/i3/p645_1.

Kaul:1983:T

[Kau83] Romesh K. Kaul. Technicolor. Reviews of Modern Physics, 55(2):449–475, April 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.449;   http://
rmp.aps.org/abstract/RMP/v55/i2/p449_1.

Keyes:1989:PDD

[Key89] Robert W. Keyes. Physics of digital devices. Reviews of Modern Physics, 61(2):279–287, April 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.279;
 http://rmp.aps.org/abstract/RMP/v61/i2/p279_1.

Kelly:1980:RPP

[KHL^{+}80] R. L. Kelly, C. P. Horne, M. J. Losty, A. Rittenberg, T. Shimada, T. G. Trippe, C. G. Wohl, G. P. Yost, N. Barash-Schmidt, C. Bricman, C. Dionisi, M. Mazzucato, L. Montanet, R. L. Crawford, M. Roos, and B. Armstrong (Particle Data Group). Review of particle properties. Reviews of Modern Physics, 52(2):S1–S286, April 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
52.S1;   http://rmp.aps.org/abstract/RMP/v52/i2/pS1_1.

Koplik:1982:SRH

[KKR82] Charles M. Koplik, Maureen F. Kaplan, and Benjamin Ross. The safety of repositories for highly radioactive wastes. Reviews of Modern Physics, 54(1):269–310, January 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.269;
 http://rmp.aps.org/abstract/RMP/v54/i1/p269_1.

Kim:1981:TER

[KLLW81] Jihn E. Kim, Paul Langacker, M. Levine, and H. H. Williams. A theoretical and experimental review of the weak neutral current: a determination of its structure and limits on deviations from the minimal SU(2)_{L} ×U(1) electroweak theory. Reviews of Modern Physics, 53(2):211–252, April 1981. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.53.211;
 http://rmp.aps.org/abstract/RMP/v53/i2/p211_1.

Kivshar:1989:DSN

[KM89] Yuri S. Kivshar and Boris A. Malomed. Dynamics of solitons in nearly integrable systems. Reviews of Modern Physics, 61 (4):763–915, October 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.763;   http:/
/rmp.aps.org/abstract/RMP/v61/i4/p763_1. See addendum

[KM91].

Kivshar:1991:ADS

[KM91] Yuri S. Kivshar and Boris A. Malomed. Addendum: Dynamics of solitons in nearly integrable systems. Reviews of Modern Physics, 63(1):211, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.63.211;   http://
rmp.aps.org/abstract/RMP/v63/i1/p211_1. See [KM89].

Kogut:1983:LGT

[Kog83] John B. Kogut. The lattice gauge theory approach to quantum chromodynamics. Reviews of Modern Physics, 55(3):775–836, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.775;   http://rmp.aps.org/abstract/
RMP/v55/i3/p775_1.

Kuo:1989:NOM

[KP89] T. K. Kuo and James Pantaleone. Neutrino oscillations in matter. Reviews of Modern Physics, 61(4):937–979, October 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
61.937;   http://rmp.aps.org/abstract/RMP/v61/i4/p937_
1.

Lane:1980:TEM

[Lan80a] N. F. Lane. The theory of electron-molecule collisions. Reviews of Modern Physics, 52(1):29–119, January 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
52.29;   http://rmp.aps.org/abstract/RMP/v52/i1/p29_1.

Langer:1980:IPF

[Lan80b] J. S. Langer. Instabilities and pattern formation in crystal growth. Reviews of Modern Physics, 52(1):1–28, January 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
52.1;   http://rmp.aps.org/abstract/RMP/v52/i1/p1_1.

Lazarus:1980:GE

[Laz80] David Lazarus. Guest editorial. Reviews of Modern Physics, 52(3):513, July 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.513;   http://
rmp.aps.org/abstract/RMP/v52/i3/p513_1.

Leggett:1987:DDT

[LCD^{+}87] A. J. Leggett, S. Chakravarty, A. T. Dorsey, Matthew P. A. Fisher, Anupam Garg, and W. Zwerger. Dynamics of the dissipative twostate system. Reviews of Modern Physics, 59(1):1–85, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.59.1;   http://rmp.aps.org/abstract/RMP/v59/

i1/p1_1. See erratum [LCD^{+}95].

Leggett:1995:EDD

[LCD^{+}95] Anthony J. Leggett, S. Chakravarty, A. T. Dorsey, Matthew P. A. Fisher, Anupam Garg, and W. Zwerger. Erratum: Dynamics of the dissipative two-state system. Reviews of Modern Physics, 67(3):725–726, July 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.67.725;   http://
rmp.aps.org/abstract/RMP/v67/i3/p725_1. See [LCD^{+}87].

Lee:1981:EXR

[LCEK81] P. A. Lee, P. H. Citrin, P. Eisenberger, and B. M. Kincaid. Extended X-ray absorption fine structure-its strengths and limitations as a structural tool. Reviews of Modern Physics, 53 (4):769–806, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.53.769;   http://
rmp.aps.org/abstract/RMP/v53/i4/p769_1.

Lederman:1989:OPP

[Led89] Leon M. Lederman. Observations in particle physics from two neutrinos to the Standard Model. Reviews of Modern Physics, 61(3):547–560, July 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.547;   http://
rmp.aps.org/abstract/RMP/v61/i3/p547_1.

Leibbrandt:1987:ING

[Lei87] George Leibbrandt. Introduction to noncovariant gauges. Reviews of Modern Physics, 59(4):1067–1119, October 1987. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.59.1067;   http://rmp.aps.org/abstract/RMP/
v59/i4/p1067_1.

Lieb:1981:TFR

[Lie81] Elliott H. Lieb. Thomas–Fermi and related theories of atoms and molecules. Reviews of Modern Physics, 53(4):603–641, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.603;   http://rmp.aps.org/abstract/

RMP/v53/i4/p603_1. See erratum [Lie82].

Lieb:1982:ETF

[Lie82] Elliott H. Lieb. Erratum: Thomas–Fermi and related theories of atoms and molecules. Reviews of Modern Physics, 54(1):311, January 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.311;   http://rmp.aps.org/abstract/
RMP/v54/i1/p311_1. See [Lie81].

Lange:1982:EEM

[LKH82a] J. Lange, Krishna Kumar, and J. H. Hamilton. E0–E2–M1 multipole admixtures of transitions in even-even nuclei. Reviews of Modern Physics, 54(1):119–194, January 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.119;
 http://rmp.aps.org/abstract/RMP/v54/i1/p119_1. See erra-

tum [LKH82b].

Lange:1982:EEE

[LKH82b] J. Lange, Krishna Kumar, and J. H. Hamilton. Erratum: E0– E2–M1 multipole admixtures of transitions in even-even nuclei. Reviews of Modern Physics, 54(2):673, April 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
54.673;   http://rmp.aps.org/abstract/RMP/v54/i2/p673_

1. See [LKH82a].

Lee:1985:DES

[LR85] Patrick A. Lee and T. V. Ramakrishnan. Disordered electronic systems. Reviews of Modern Physics, 57(2):287–337, April 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.287;   http://rmp.aps.org/abstract/
RMP/v57/i2/p287_1.

Mackey:1989:DOI

[Mac89] Michael C. Mackey. The dynamic origin of increasing entropy. Reviews of Modern Physics, 61(4):981–1015, October 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
61.981;   http://rmp.aps.org/abstract/RMP/v61/i4/p981_
1.

Martin:1988:SRC

[Mar88] Ph. A. Martin. Sum rules in charged fluids. Reviews of Modern Physics, 60(4):1075–1127, October 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.60.1075;   http:/
/rmp.aps.org/abstract/RMP/v60/i4/p1075_1.

Mattis:1986:FBP

[Mat86] Daniel C. Mattis. The few-body problem on a lattice. Reviews of Modern Physics, 58(2):361–379, April 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.361;
 http://rmp.aps.org/abstract/RMP/v58/i2/p361_1.

McLerran:1986:PQG

[McL86] Larry McLerran. The physics of the quark-gluon plasma. Reviews of Modern Physics, 58(4):1021–1064, October 1986. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.58.1021;   http://rmp.aps.org/abstract/RMP/
v58/i4/p1021_1.

Michel:1980:SDB

[Mic80] Louis Michel. Symmetry defects and broken symmetry. Configurations hidden symmetry. Reviews of Modern Physics, 52(3):617–651, July 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.617;   http://rmp.aps.org/abstract/
RMP/v52/i3/p617_1.

Michel:1982:TPM

[Mic82] F. Curtis Michel. Theory of pulsar magnetospheres. Reviews of Modern Physics, 54(1):1–66, January 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.1;
http://rmp.aps.org/abstract/RMP/v54/i1/p1_1.

Minnhagen:1987:TDC

[Min87] Petter Minnhagen. The two-dimensional Coulomb gas, vortex unbinding, and superfluid-superconducting films. Reviews of Modern Physics, 59(4):1001–1066, October 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.59.1001;   http:/
/rmp.aps.org/abstract/RMP/v59/i4/p1001_1.

Mohring:1983:PFA

[MMO83] Willi Möhring, Ernst-August Müller, and Frank Obermeier. Problems in flow acoustics. Reviews of Modern Physics, 55(3):707–724, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.707;   http://rmp.aps.org/abstract/
RMP/v55/i3/p707_1.

Montvay:1987:NCH

[Mon87] I. Montvay. Numerical calculation of hadron masses in quantum chromodynamics. Reviews of Modern Physics, 59(1):263–285, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.263;   http://rmp.aps.org/abstract/
RMP/v59/i1/p263_1.

Moskovits:1985:SES

[Mos85] Martin Moskovits. Surface-enhanced spectroscopy. Reviews of Modern Physics, 57(3):783–826, July 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.783;
 http://rmp.aps.org/abstract/RMP/v57/i3/p783_1.

Menikoff:1989:RPF

[MP89] Ralph Menikoff and Bradley J. Plohr. The Riemann problem for fluid flow of real materials. Reviews of Modern Physics, 61 (1):75–130, January 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.61.75;   http://rmp.aps.
org/abstract/RMP/v61/i1/p75_1.

Mints:1981:CSS

[MR81] R. G. Mints and A. L. Rakhmanov. Critical state stability in type-II superconductors and superconducting-normal-metal composites. Reviews of Modern Physics, 53(3):551–592, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
53.551;   http://rmp.aps.org/abstract/RMP/v53/i3/p551_
1.

Manson:1982:PAD

[MS82] Steven T. Manson and Anthony F. Starace. Photoelectron angular distributions: energy dependence for s subshells. Reviews of Modern Physics, 54(2):389–405, April 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.389;
 http://rmp.aps.org/abstract/RMP/v54/i2/p389_1.

Myhrer:1988:NNF

[MW88] Fred Myhrer and J. Wroldsen. The nucleon-nucleon force and the quark degrees of freedom. Reviews of Modern Physics, 60(3):629–661, July 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.60.629;   http://
rmp.aps.org/abstract/RMP/v60/i3/p629_1.

Negele:1982:MFT

[Neg82] J. W. Negele. The mean-field theory of nuclear structure and dynamics. Reviews of Modern Physics, 54(4):913–1015, October 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.913;   http://rmp.aps.org/abstract/
RMP/v54/i4/p913_1.

ODonnell:1981:RDM

[O’D81] Patrick J. O’Donnell. Radiative decays of mesons. Reviews of Modern Physics, 53(4):673–685, October 1981. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.53.673;
 http://rmp.aps.org/abstract/RMP/v53/i4/p673_1.

Ostriker:1988:AB

[OM88] Jeremiah P. Ostriker and Christopher F. McKee. Astrophysical blastwaves. Reviews of Modern Physics, 60(1):1–68, January 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.60.1;   http://rmp.aps.org/abstract/RMP/v60/
i1/p1_1.

Olariu:1985:QEE

[OP85] S. Olariu and I. Iovitzu Popescu. The quantum effects of electromagnetic fluxes. Reviews of Modern Physics, 57(2):339–436, April 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.339;   http://rmp.aps.org/abstract/
RMP/v57/i2/p339_1.

Ott:1981:SAC

[Ott81] Edward Ott. Strange attractors and chaotic motions of dynamical systems. Reviews of Modern Physics, 53(4):655–671, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.655;   http://rmp.aps.org/abstract/
RMP/v53/i4/p655_1.

Owens:1987:LMT

[Owe87] J. F. Owens. Large-momentum-transfer production of direct photons, jets, and particles. Reviews of Modern Physics, 59 (2):465–503, April 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.59.465;   http://
rmp.aps.org/abstract/RMP/v59/i2/p465_1.

Patterson:1988:MSS

[Pat88] Bruce D. Patterson. Muonium states in semiconductors. Reviews of Modern Physics, 60(1):69–159, January 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
60.69;   http://rmp.aps.org/abstract/RMP/v60/i1/p69_1.

Paul:1982:PA

[Pau82] H. Paul. Photon antibunching. Reviews of Modern Physics, 54 (4):1061–1102, October 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.54.1061;   http://
rmp.aps.org/abstract/RMP/v54/i4/p1061_1.

Paul:1986:IBI

[Pau86] H. Paul. Interference between independent photons. Reviews of Modern Physics, 58(1):209–231, January 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.209;
 http://rmp.aps.org/abstract/RMP/v58/i1/p209_1.

Pickett:1989:ESH

[Pic89a] Warren E. Pickett. Electronic structure of the high-temperature oxide superconductors. Reviews of Modern Physics, 61(2):433–512, April 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.61.433;   http://rmp.aps.org/abstract/

RMP/v61/i2/p433_1. See erratum [Pic89b].

Pickett:1989:EES

[Pic89b] Warren E. Pickett. Erratum: Electronic structure of the high-temperature oxide superconductors. Reviews of Modern Physics, 61(3):749, July 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.749;   http://
rmp.aps.org/abstract/RMP/v61/i3/p749_1. See [Pic89a].

Pines:1986:ECI

[Pin86] David Pines. Editorial: Color illustrations in Reviews of Modern Physics. Reviews of Modern Physics, 58(3):517, July 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
58.517;   http://rmp.aps.org/abstract/RMP/v58/i3/p517_

Pines:1987:ECC

[Pin87] David Pines. Editorial: Coverage of complex systems. Reviews of Modern Physics, 59(3):531, July 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.531;
 http://rmp.aps.org/abstract/RMP/v59/i3/p531_1.

Peixoto:1984:PC

[PO84] José P. Peixóto and Abraham H. Oort. Physics of climate. Reviews of Modern Physics, 56(3):365–429, July 1984. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.56.365;
 http://rmp.aps.org/abstract/RMP/v56/i3/p365_1.

Patel:1981:POS

[PT81] C. K. N. Patel and A. C. Tam. Pulsed optoacoustic spectroscopy of condensed matter. Reviews of Modern Physics, 53(3):517–550, July 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.517;   http://rmp.aps.org/abstract/
RMP/v53/i3/p517_1.

Roger:1983:MS

[RHD83] M. Roger, J. H. Hetherington, and J. M. Delrieu. Magnetism in solid ^{3}He. Reviews of Modern Physics, 55(1):1–64, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.55.1;   http://rmp.aps.org/abstract/RMP/v55/
i1/p1_1.

Rich:1981:REA

[Ric81] Arthur Rich. Recent experimental advances in positronium research. Reviews of Modern Physics, 53(1):127–165, January 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.127;   http://rmp.aps.org/abstract/
RMP/v53/i1/p127_1.

Rigden:1986:QSP

[Rig86] John S. Rigden. Quantum states and precession: The two discoveries of NMR. Reviews of Modern Physics, 58(2):433–448, April 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.433;   http://rmp.aps.org/abstract/
RMP/v58/i2/p433_1.

Rudd:1985:EPP

[RKMG85] M. E. Rudd, Y.-K. Kim, D. H. Madison, and J. W. Gallagher. Electron production in proton collisions: total cross sections. Reviews of Modern Physics, 57(4):965–994, October 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.57.965;
 http://rmp.aps.org/abstract/RMP/v57/i4/p965_1.

Rammer:1986:QFT

[RS86] J. Rammer and H. Smith. Quantum field-theoretical methods in transport theory of metals. Reviews of Modern Physics, 58(2):323–359, April 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.58.323;   http://
rmp.aps.org/abstract/RMP/v58/i2/p323_1.

Rammal:1986:UP

[RTV86] R. Rammal, G. Toulouse, and M. A. Virasoro. Ultrametricity for physicists. Reviews of Modern Physics, 58(3):765–788, July 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.765;   http://rmp.aps.org/abstract/
RMP/v58/i3/p765_1.

Rubbia:1985:EOI

[Rub85] Carlo Rubbia. Experimental observation of the intermediate vector bosons W^{+}, W^{−}, and Z^{0}. Reviews of Modern Physics, 57(3):699–722, July 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.57.699;   http://
rmp.aps.org/abstract/RMP/v57/i3/p699_1.

Ruska:1987:DEM

[Rus87] Ernst Ruska. The development of the electron microscope and of electron microscopy. Reviews of Modern Physics, 59(3):627–638, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.627;   http://rmp.aps.org/abstract/
RMP/v59/i3/p627_1.

Salam:1980:GUF

[Sal80] Abdus Salam. Gauge unification of fundamental forces. Reviews of Modern Physics, 52(3):525–538, July 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.525;
 http://rmp.aps.org/abstract/RMP/v52/i3/p525_1. See com-

ments and corrections [Ynd81].

Sarazin:1986:XRE

[Sar86] Craig L. Sarazin. X-ray emission from clusters of galaxies. Reviews of Modern Physics, 58(1):1–115, January 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.1;
http://rmp.aps.org/abstract/RMP/v58/i1/p1_1.

Savit:1980:DFT

[Sav80] Robert Savit. Duality in field theory and statistical systems. Reviews of Modern Physics, 52(2):453–487, April 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.453;
 http://rmp.aps.org/abstract/RMP/v52/i2/p453_1.

Schawlow:1982:SNL

[Sch82] Arthur L. Schawlow. Spectroscopy in a new light. Reviews of Modern Physics, 54(3):697–707, July 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.697;
 http://rmp.aps.org/abstract/RMP/v54/i3/p697_1.

Schweber:1986:FVS

[Sch86] Silvan S. Schweber. Feynman and the visualization of spacetime processes. Reviews of Modern Physics, 58(2):449–508, April 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.449;   http://rmp.aps.org/abstract/
RMP/v58/i2/p449_1.

Schatz:1989:ARE

[Sch89a] George C. Schatz. The analytical representation of electronic potential-energy surfaces. Reviews of Modern Physics, 61(3):669–688, July 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.669;   http://
rmp.aps.org/abstract/RMP/v61/i3/p669_1.

Schwartz:1989:FHE

[Sch89b] Mel Schwartz. The first high-energy neutrino experiment. Reviews of Modern Physics, 61(3):527–532, July 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.527;
 http://rmp.aps.org/abstract/RMP/v61/i3/p527_1.

Shimizu:1989:PFR

[SGB^{+}89] Y. R. Shimizu, J. D. Garrett, R. A. Broglia, M. Gallardo, and E. Vigezzi. Pairing fluctuations in rapidly rotating nuclei. Reviews of Modern Physics, 61(1):131–168, January 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.131;
 http://rmp.aps.org/abstract/RMP/v61/i1/p131_1.

Siegbahn:1982:ESA

[Sie82] Kai Siegbahn. Electron spectroscopy for atoms, molecules, and condensed matter. Reviews of Modern Physics, 54(3):709–728, July 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.709;   http://rmp.aps.org/abstract/
RMP/v54/i3/p709_1.

Silvera:1980:SMH

[Sil80] Isaac F. Silvera. The solid molecular hydrogens in the condensed phase: Fundamentals and static properties. Reviews of Modern Physics, 52(2):393–452, April 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.393;
 http://rmp.aps.org/abstract/RMP/v52/i2/p393_1.

Schultz:1988:IPB

[SL88] Peter J. Schultz and K. G. Lynn. Interaction of positron beams with surfaces, thin films, and interfaces. Reviews of Modern Physics, 60(3):701–779, July 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.701;
 http://rmp.aps.org/abstract/RMP/v60/i3/p701_1.

Sonin:1987:VOH

[Son87] E. B. Sonin. Vortex oscillations and hydrodynamics of rotating superfluids. Reviews of Modern Physics, 59(1):87–155, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.59.87;   http://rmp.aps.org/
abstract/RMP/v59/i1/p87_1.

Swartz:1989:TBR

[SP89] E. T. Swartz and R. O. Pohl. Thermal boundary resistance. Reviews of Modern Physics, 61(3):605–668, July 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.605;
 http://rmp.aps.org/abstract/RMP/v61/i3/p605_1.

Spohn:1980:KEH

[Spo80] Herbert Spohn. Kinetic equations from Hamiltonian dynamics: Markovian limits. Reviews of Modern Physics, 52(3):569–615, July 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.569;   http://rmp.aps.org/abstract/
RMP/v52/i3/p569_1.

Stewart:1984:HFS

[Ste84] G. R. Stewart. Heavy-fermion systems. Reviews of Modern Physics, 56(4):755–787, October 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.56.755;   http://
rmp.aps.org/abstract/RMP/v56/i4/p755_1.

Stenholm:1986:STL

[Ste86] Stig Stenholm. The semiclassical theory of laser cooling. Reviews of Modern Physics, 58(3):699–739, July 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.699;
 http://rmp.aps.org/abstract/RMP/v58/i3/p699_1.

Steinberger:1989:EHE

[Ste89] J. Steinberger. Experiments with high-energy neutrino beams. Reviews of Modern Physics, 61(3):533–545, July 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.533;
 http://rmp.aps.org/abstract/RMP/v61/i3/p533_1.

Stacey:1987:GLG

[STM^{+}87] F. D. Stacey, G. J. Tuck, G. I. Moore, S. C. Holding, B. D. Goodwin, and R. Zhou. Geophysics and the law of gravity. Reviews of Modern Physics, 59(1):157–174, January 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.157;
 http://rmp.aps.org/abstract/RMP/v59/i1/p157_1.

Strandburg:1988:TDM

[Str88] Katherine J. Strandburg. Two-dimensional melting. Reviews of Modern Physics, 60(1):161–207, January 1988. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.60.161;
 http://rmp.aps.org/abstract/RMP/v60/i1/p161_1. See erra-

tum [Str89].

Strandburg:1989:ETD

[Str89] Katherine J. Strandburg. Erratum: Two-dimensional melting. Reviews of Modern Physics, 61(3):747, July 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
61.747;   http://rmp.aps.org/abstract/RMP/v61/i3/p747_

1. See [Str88].

Salomaa:1987:QVS

[SV87] M. M. Salomaa and G. E. Volovik. Quantized vortices in superfluid ^{3}He. Reviews of Modern Physics, 59(3):533–613, July 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.533;   http://rmp.aps.org/abstract/

RMP/v59/i3/p533_1. See erratum [SV88].

Salomaa:1988:EQV

[SV88] M. M. Salomaa and G. E. Volovik. Erratum: Quantized vortices in superfluid ^{3}He. Reviews of Modern Physics, 60(2):573, April 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.573;   http://rmp.aps.org/abstract/
RMP/v60/i2/p573_1. See [SV87].

Shandarin:1989:LSS

[SZ89] S. F. Shandarin and Ya. B. Zeldovich. The large-scale structure of the universe: Turbulence, intermittency, structures in a selfgravitating medium. Reviews of Modern Physics, 61(2):185–220, April 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.61.185;   http://rmp.aps.org/abstract/
RMP/v61/i2/p185_1.

Tam:1986:APS

[Tam86] Andrew C. Tam. Applications of photoacoustic sensing techniques. Reviews of Modern Physics, 58(2):381–431, April 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
58.381;   http://rmp.aps.org/abstract/RMP/v58/i2/p381_
1.

Taylor:1986:RMR

[Tay86] J. B. Taylor. Relaxation and magnetic reconnection in plasmas. Reviews of Modern Physics, 58(3):741–763, July 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.741;
 http://rmp.aps.org/abstract/RMP/v58/i3/p741_1.

Tucker:1985:QDM

[TF85] John R. Tucker and Marc J. Feldman. Quantum detection at millimeter wavelengths. Reviews of Modern Physics, 57(4): 1055–1113, October 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.57.1055;   http://
rmp.aps.org/abstract/RMP/v57/i4/p1055_1.

Thacker:1981:EIQ

[Tha81] H. B. Thacker. Exact integrability in quantum field theory and statistical systems. Reviews of Modern Physics, 53(2):253–285, April 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.253;   http://rmp.aps.org/abstract/
RMP/v53/i2/p253_1.

Thorne:1980:GWR

[Tho80a] Kip S. Thorne. Gravitational-wave research: Current status and future prospects. Reviews of Modern Physics, 52(2):285–297, April 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.52.285;   http://rmp.aps.org/abstract/
RMP/v52/i2/p285_1.

Thorne:1980:MEG

[Tho80b] Kip S. Thorne. Multipole expansions of gravitational radiation. Reviews of Modern Physics, 52(2):299–339, April 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.299;
 http://rmp.aps.org/abstract/RMP/v52/i2/p299_1.

Tonomura:1987:AEH

[Ton87] Akira Tonomura. Applications of electron holography. Reviews of Modern Physics, 59(3):639–669, July 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.639;
 http://rmp.aps.org/abstract/RMP/v59/i3/p639_1.

Trimble:1982:SPE

[Tri82] Virginia Trimble. Supernovae. Part I: the events. Reviews of Modern Physics, 54(4):1183–1224, October 1982. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.54.1183;   http://rmp.aps.org/abstract/RMP/
v54/i4/p1183_1.

Trimble:1983:SPI

[Tri83] Virginia Trimble. Supernovae. Part II: the aftermath. Reviews of Modern Physics, 55(2):511–563, April 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.511;
 http://rmp.aps.org/abstract/RMP/v55/i2/p511_1.

Trimble:1988:GSS

[Tri88] Virginia Trimble. 1987A: The greatest supernova since Kepler. Reviews of Modern Physics, 60(4):859–871, October 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
60.859;   http://rmp.aps.org/abstract/RMP/v60/i4/p859_

vanBeijeren:1982:TPS

[vB82] Henk van Beijeren. Transport properties of stochastic Lorentz models. Reviews of Modern Physics, 54(1):195–234, January 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.54.195;   http://rmp.aps.org/abstract/
RMP/v54/i1/p195_1.

vanderMeer:1985:SCA

[vdM85] S. van der Meer. Stochastic cooling and the accumulation of antiprotons. Reviews of Modern Physics, 57(3):689–697, July 1985. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.57.689;   http://rmp.aps.org/abstract/
RMP/v57/i3/p689_1.

Vivaldi:1984:WIM

[Viv84] Franco Vivaldi. Weak instabilities in many-dimensional Hamiltonian systems. Reviews of Modern Physics, 56(4):737–754, October 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.56.737;   http://rmp.aps.org/abstract/
RMP/v56/i4/p737_1.

vonKlitzing:1986:QHE

[vK86] Klaus von Klitzing. The quantized Hall effect. Reviews of Modern Physics, 58(3):519–531, July 1986. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.58.519;
 http://rmp.aps.org/abstract/RMP/v58/i3/p519_1.

Vollhardt:1984:NAL

[Vol84] Dieter Vollhardt. Normal ^{3}He: an almost localized Fermi liquid. Reviews of Modern Physics, 56(1):99–120, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
56.99;   http://rmp.aps.org/abstract/RMP/v56/i1/p99_1.

Wilson:1985:RAP

[WAA^{+}85] Richard Wilson, Kamal J. Araj, Augustine O. Allen, Peter Auer, David G. Boulware, Fred Finlayson, Simon Goren, Clark Ice, Leon Lidofsky, Allen Lee Sessoms, Mary L. Shoaf, Irving Spiewak, Thomas Tombrello, Herbert S. Gutowsky, Andreas Acrivos, Herman Feshbach, and William A. Fowler (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society of the study group on radionuclide release from severe accidents at nuclear power plants. Reviews of Modern Physics, 57(3):S1–S144, July 1985. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
57.S1;   http://rmp.aps.org/abstract/RMP/v57/i3/pS1_1.

Wohl:1984:RPP

[WCR^{+}84] C. G. Wohl, R. N. Cahn, A. Rittenberg, T. G. Trippe, G. P. Yost, F. C. Porter, J. J. Hernandez, L. Montanet, R. E. Hendrick, R. L. Crawford, M. Roos, N. A. Törnqvist, G. Höhler, M. Aguilar-Benitez, T. Shimada, M. J. Losty, G. P. Gopal, Ch. Walck, R. E. Shrock, R. Frosch, L. D. Roper, W. P. Trower, and B. Armstrong (Particle Data Group). Review of particle properties. Reviews of Modern Physics, 56(2):S1–S299, April 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
56.S1;   http://rmp.aps.org/abstract/RMP/v56/i2/pS1_1.

Weinberg:1980:CFU

[Wei80] Steven Weinberg. Conceptual foundations of the unified theory of weak and electromagnetic interactions. Reviews of Modern Physics, 52(3):515–523, July 1980. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.52.515;
 http://rmp.aps.org/abstract/RMP/v52/i3/p515_1.

Weissman:1988:NOS

[Wei88] M. B. Weissman. 1/f noise and other slow, nonexponential kinetics in condensed matter. Reviews of Modern Physics, 60(2):537–571, April 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.537;   http://rmp.aps.org/abstract/
RMP/v60/i2/p537_1.

Weinberg:1989:CCP

[Wei89] Steven Weinberg. The cosmological constant problem. Reviews of Modern Physics, 61(1):1–23, January 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.1;
http://rmp.aps.org/abstract/RMP/v61/i1/p1_1.

Whitley:1984:RGCa

[Whi84a] Stanley Whitley. Review of the gas centrifuge until 1962. Part I: Principles of separation physics. Reviews of Modern Physics, 56 (1):41–66, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.56.41;   http://rmp.aps.
org/abstract/RMP/v56/i1/p41_1.

Whitley:1984:RGCb

[Whi84b] Stanley Whitley. Review of the gas centrifuge until 1962. Part II: Principles of high-speed rotation. Reviews of Modern Physics, 56(1):67–97, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.56.67;   http://rmp.aps.
org/abstract/RMP/v56/i1/p67_1.

White:1986:RR

[Whi86] R. B. White. Resistive reconnection. Reviews of Modern Physics, 58(1):183–207, January 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.58.183;   http://
rmp.aps.org/abstract/RMP/v58/i1/p183_1.

Wilson:1983:RGC

[Wil83] Kenneth G. Wilson. The renormalization group and critical phenomena. Reviews of Modern Physics, 55(3):583–600, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.55.583;   http://rmp.aps.org/abstract/
RMP/v55/i3/p583_1.

Wright:1989:CLB

[WM89] David C. Wright and N. David Mermin. Crystalline liquids: the blue phases. Reviews of Modern Physics, 61(2):385–432, April 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.61.385;   http://rmp.aps.org/abstract/
RMP/v61/i2/p385_1.

Wolfram:1983:SMC

[Wol83] Stephen Wolfram. Statistical mechanics of cellular automata. Reviews of Modern Physics, 55(3):601–644, July 1983. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.55.601;
 http://rmp.aps.org/abstract/RMP/v55/i3/p601_1.

Weller:1980:CRP

[WR80] H. R. Weller and N. R. Roberson. Capture reactions with protons, neutrons, and alpha particles. Reviews of Modern Physics, 52(4):699–724, October 1980. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.52.699;   http://
rmp.aps.org/abstract/RMP/v52/i4/p699_1.

Wu:1982:PM

[Wu82] F. Y. Wu. The Potts model. Reviews of Modern Physics, 54 (1):235–268, January 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.54.235;   http:/
/rmp.aps.org/abstract/RMP/v54/i1/p235_1.  See erratum

[Wu83].

Wu:1983:EPM

[Wu83] F. Y. Wu. Erratum: The Potts model. Reviews of Modern Physics, 55(1):315, January 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.55.315;   http://
rmp.aps.org/abstract/RMP/v55/i1/p315_1. See [Wu82].

Yaffe:1982:LLC

[Yaf82] Laurence G. Yaffe. Large N limits as classical mechanics. Reviews of Modern Physics, 54(2):407–435, April 1982. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.54.407;
 http://rmp.aps.org/abstract/RMP/v54/i2/p407_1.

Yennie:1987:IQH

[Yen87] D. R. Yennie. Integral quantum Hall effect for nonspecialists. Reviews of Modern Physics, 59(3):781–824, July 1987. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.59.781;
 http://rmp.aps.org/abstract/RMP/v59/i3/p781_1.

Yamamoto:1986:PMI

[YH86] Y. Yamamoto and H. A. Haus. Preparation, measurement and information capacity of optical quantum states. Reviews of Modern Physics, 58(4):1001–1020, October 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.58.1001;   http:/
/rmp.aps.org/abstract/RMP/v58/i4/p1001_1.

Yndurain:1981:LE

[Ynd81] F. J. Ynduráin. A letter to the editor. Reviews of Modern Physics, 53(2):359, April 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.53.359;   http://
rmp.aps.org/abstract/RMP/v53/i2/p359_1. Comments and cor-

rections to [Sal80].

Zeilinger:1988:SDS

[ZGS^{+}88] Anton Zeilinger, Roland Gähler, C. G. Shull, Wolfgang Treimer, and Walter Mampe. Single- and double-slit diffraction of neutrons. Reviews of Modern Physics, 60(4):1067–1073, October 1988. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.60.1067;   http://rmp.aps.org/abstract/
RMP/v60/i4/p1067_1.